OpenStudy (anonymous):

I've been trying to figure this out for a long time. Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values.

OpenStudy (anonymous):

Solve for x by plugging in y: x - 62,500 = 0.25(125,000) x - 62,500 = 31,250 x = 93,750

OpenStudy (anonymous):

how do you do that? I'm really confused. Can you do it for me?

OpenStudy (anonymous):

Solve for x by plugging in y: x - 62,500 = 0.25(125,000) (multiply the total amount of voters by the percent sally received) x - 62,500 = 31,250 (add 62,500 to each side to get x alone) x = 93,750 = Number of votes Maya received

OpenStudy (anonymous):

so the final answer is 93,750

OpenStudy (anonymous):

thats how many votes maya received yes

OpenStudy (anonymous):

but since you aren't supposed to put place holders it'd be 93750